OpenStudy (anonymous):

What is the equation of the line that passes through (–2, 3) and is parallel to 2x + 3y = 6?

OpenStudy (anonymous):

The slope of the parallel line will be the same as the slope from the given line. if you put the given line into y=mx+b form, you can read the slope, m. 2x-3y=6 2x-6=3y y=(2/3)x-2 So the slope of the line is m=2/3. Now you have a point and the slope, so you can use the point-slope formula to find the line: y-1=(2/3)(x+3) y=(2/3)x+2+1 or y=(2/3)x+3

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

\[ y= -\frac{2x}{3}+ \frac{5}{3} \\ \]

OpenStudy (anonymous):

I hoped I helped you. c: